OpenStudy (anonymous):

Find the equation in standard form which passes through (–1, 4) and (0, 9).

OpenStudy (anonymous):

slope m=y2-y1/x2-x1=(9-4)/(0+1)=5 y-y1=m(x-x1) y-9=5(x) y=5x+9

OpenStudy (zepp):

What is the slope of this line?

OpenStudy (anonymous):

equation in standard form which passes through (–1, 4) and (0, 9). is y=5x+9

OpenStudy (anonymous):

Your my sleek-feathered one!

OpenStudy (anonymous):

lmfao

OpenStudy (zepp):

@sai1234 Please show steps when posting your answer, giving it out directly is against the Code of Conduct. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

i alredy given steps above

OpenStudy (zepp):

I see, my bad.

OpenStudy (anonymous):

@zepp he is right

OpenStudy (anonymous):

I have received help, i dont even need the answer. Just help